Extended Constructed Response
a drawing of a wall that is to be covered with either wallpaper or paint. The wall is 6 ft. high and 20 ft. wide. The window, mirror, and fireplace are not to be painted or papered. The window measures 18 in. wide and 14 ft. high. The fireplace is 6 ft. wide and 4 ft. high, while the mirror above the fireplace is 6 ft. wide and 3 ft. high.
Pay attention to the units! You need to make them all the same.
How many square feet of wallpaper are needed to cover the wall?
Explain how you arrived at your answer.

Answer by mananth(16946)

21.
Wall paper required = Area of wall - ( area of window + area of fireplace + area of mirror)
Area of wall = 6*20 = 120 ft^2


Area of window = 14*1.5 =21 ft^2
Area of fire place 6*4 = 24 ft^2
Area of mirror = 6*3 = 18 ft^2
--------
63 ft^2
Area to be painted / wall papered = 120 -63 =57ft^2